OpenStudy (anonymous):

find the x and y intercepts 4x-7y=-28

OpenStudy (a_clan):

put x=0 in this eqn , to get y-intercept

OpenStudy (a_clan):

put y=0 in this eqn , to get x-intercept

OpenStudy (anonymous):

x intercept y=0 4x-7(0)=-28 x=-28/4=-7 y intercepts x=0 4(0)-7y=-28 y-28/7=4